# Temporally Consistent Amodal Completion for 3D Human-Object Interaction Reconstruction

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–30, 2025; Dublin, Irelandsubmissionid: 532ccs: Computing methodologies Shape inferenceccs: Computing methodologies Reconstruction
@@ 1. Introduction
Amodal completion (Zhou et al., 2023b; Chen et al., 2023; Author et al., 2023; Li et al., 2022; Nguyen et al., 2022; Wu et al., 2022; Author and Author, 2023) has emerged as a promising strategy to mitigate these occlusion challenges by inferring the complete structure of partially hidden regions. However, most existing approaches are designed for static scenes. Although they perform well in controlled environments, they often falter in dynamic, real-world settings. A key limitation is their reliance on image-level completion, which overlooks the rich temporal context provided by adjacent sequences in HOI cases. Without leveraging temporal information, reconstructions derived from amodal-completed images can suffer from inconsistencies, ultimately compromising the overall quality and stability of the models.
@@ 2.2. Amodal Completion
Amodal completion focuses on inferring the hidden shapes and appearances of occluded objects. Recently, various works leverage generative models to reconstruct missing regions. For instance, pix2gestalt (Author et al., 2023) synthesizes “wholes” from partial observations, while Progressive Mixed Context Diffusion (Author and Author, 2023) expands occluded areas with context-awareness. Other studies propose transformer-based amodal segmentation (Zhou et al., 2023b), leverage 3D shape priors to guide amodal masks (Chen et al., 2023), or learn compositional scene representations to handle complex occluders (Li et al., 2022; Nguyen et al., 2022). Self-supervised frameworks alsowemerged for amodal completion by aligning visible parts with plausible hidden geometry (Wu et al., 2022; Kim et al., 2023). However, these single-image methods often struggle with dynamic human-object interactions, where severe occlusions may shift from frame to frame—leading to temporally inconsistent reconstructions. As an alternatives, video inpainting methods (Lu et al., 2023) has also been introduced, however these methods are usually for removing subjects well fitted to background and environment, showing limitation on HOI cases. Our approach addresses these issues by integrating temporal cues, ensuring stable amodal completions across sequences.
@@ 3. Method
3.1. Problem Formulation
We address the problem of filling in occluded regions in a sequence of video frames such that the completed regions are both semantically plausible and temporally coherent. For each frame, let $I_{in}\in\mathbb{R}^{H\times W\times 3}$ be the segmented image containing only the visible regions, and let $M_{in}\in\{0,1\}^{H\times W}$ be a binary mask, where 1 indicates the occluded regions to be inpainted. In addition, we provide a temporally-aware latent feature $z_{in}$, aggregated from temporally aligned features across neighboring frames, and a text prompt $P$ as textual guidance. We formalize the inpainting process with the diffusion model (Rombach et al., 2022) guided by both latent and mask inputs as:
@@ 3.2. Bidirectional Temporal Feature Warping
(4) $\hat{z}^{t}_{\tau}=\mathrm{Warp}(z^{\tau},o^{t}_{\tau}),\quad\forall\,\tau\in\mathcal{S}(t),$
@@ 3.4. Template-free Occlusion Identification
(7) $M^{projected}_{occludee}=\text{Hull}(C).$
@@
(9) $M_{occlusion}=M_{union}\cap M_{occluder}.$
With this approach, we precisely identify the occluded region of the target object without relying on any predefined templates, significantly enhancing the generalizability and robustness of our approach.
@@ 3.6. Application in 3D Reconstruction
(12) $\mathcal{L}_{obj}=\sum_{t}\mathcal{L}_{photo}\left(R(\mathcal{G}_{obj},H_{t}\circ X_{t}),\;\tilde{I}_{out}^{t}\right),$
@@ 4.3. Temporally Consistent Amodal Completion
We quantitatively evaluate our approach against several state-of-the-art baselines, including Pix2gestalt (Author et al., 2023), Stable Diffusion Inpainting (Rombach et al., 2022), LaMa (Suvorov et al., 2022), and VDT (Lu et al., 2023), initialized with random noise. The results are summarized in Table 1, covering both human and object categories, as the human is treated as one of 19 object classes in our experiments, and only frames with an object occlusion ratio between 15% and 70% are considered. Our method consistently achieves the best performance across amodal completion metrics, IoU and CLIP, on both BEHAVE and InterCap. This indicates a higher fidelity in reconstructing occluded regions, validating the effectiveness of our template-free regioning strategy in complex HOI cases. To assess ideal condition performance, we report results using ground-truth masks in the last row.
@@
Figure 5. Qualitative results of 3D Reconstruction: Training on our temporally consistent completions (left) produces clearer geometry and fewer artifacts than single-frame (middle) or no completion (right).
@@
Table 2. Quantitative results on 3D reconstruction from the BEHAVE dataset (Bhatnagar et al., 2022): We report PSNR-M, SSIM-M, and LPIPS-M, computed within tight bounding boxes around humans or objects. Our method outperforms all baselines across metrics.
@@ C.1. Baselines
Table 5. 3D bounding box size and volume for each object in BEHAVE (Bhatnagar et al., 2022), sorted by descending size. boxtiny and toolbox have the smallest spatial extent and are excluded from 3D reconstruction experiments.
@@ D. Evaluation
To evaluate the quality of amodal completion, we use two complementary metrics: CLIP Score (Radford et al., 2021) and Intersection over Union (IoU). The CLIP score quantifies the semantic alignment between each inpainted image and a corresponding category-level textual prompt using CLIP’s vision–language embedding space. This allows us to assess whether the completed region semantically resembles the intended object class. To reduce the impact of unrelated background pixels, we compute CLIP scores within tight bounding boxes centered on the reconstructed objects.

### v2

# Occlusion-Aware Temporally Consistent Amodal Completion for 3D Human-Object Interaction Reconstruction

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754769isbn: 979-8-4007-2035-2/2025/10submissionid: 532ccs: Computing methodologies Shape inferenceccs: Computing methodologies Reconstruction
@@ 1. Introduction
Amodal completion (Zhou et al., 2023b; Chen et al., 2023; Ozguroglu et al., 2024; Li et al., 2022; Nguyen et al., 2022; Wu et al., 2022; Xu et al., 2024; Chi et al., 2025) has emerged as a promising strategy to mitigate these occlusion challenges by inferring the complete structure of partially hidden regions. However, most existing approaches are designed for static scenes. Although they perform well in controlled environments, they often falter in dynamic, real-world settings. A key limitation is their reliance on image-level completion, which overlooks the rich temporal context provided by adjacent sequences in HOI cases. Without leveraging temporal information, reconstructions derived from amodal-completed images can suffer from inconsistencies, ultimately compromising the overall quality and stability of the models.
@@ 2.2. Amodal Completion
Amodal completion focuses on inferring the hidden shapes and appearances of occluded objects. Recently, various works leverage generative models to reconstruct missing regions. For instance, Chi et al. (Chi et al., 2025) completes occlusion using contact points while pix2gestalt (Ozguroglu et al., 2024) synthesizes “wholes” from partial observations and Xu et al. (Xu et al., 2024) expands occluded areas with context-awareness. Other studies propose transformer-based amodal segmentation (Zhou et al., 2023b), leverage 3D shape priors to guide amodal masks (Chen et al., 2023), or learn compositional scene representations to handle complex occluders (Li et al., 2022; Nguyen et al., 2022). Self-supervised frameworks alsowemerged for amodal completion by aligning visible parts with plausible hidden geometry (Wu et al., 2022; Kim et al., 2023). However, these single-image methods often struggle with dynamic human-object interactions, where severe occlusions may shift from frame to frame—leading to temporally inconsistent reconstructions. As an alternatives, video inpainting methods (Lu et al., 2023) has also been introduced, however these methods are usually for removing subjects well fitted to background and environment, showing limitation on HOI cases. Our approach addresses these issues by integrating temporal cues, ensuring stable amodal completions across sequences.
@@ 3. Method
3.1. Problem Definition
We address the problem of filling in occluded regions in a sequence of video frames such that the completed regions are both semantically plausible and temporally coherent. For each frame, let $I_{in}\in\mathbb{R}^{H\times W\times 3}$ be the segmented image containing only the visible regions, and let $M_{in}\in\{0,1\}^{H\times W}$ be a binary mask, where 1 indicates the occluded regions to be inpainted. In addition, we provide a temporally-aware latent feature $z_{in}$, aggregated from temporally aligned features across neighboring frames, and a text prompt $P$ as textual guidance. Adapting the problem definition of (Xu et al., 2024), we formalize the inpainting process with the diffusion model (Rombach et al., 2022) guided by both latent and mask inputs as:
@@ 3.2. Bidirectional Temporal Feature Warping
(4) $\hat{z}^{t}_{\tau}=\mathrm{Warp}(z^{\tau},o^{t}_{\tau}),\quad\forall\,\tau\in\mathcal{S}(t).$
@@ 3.4. Template-free Occlusion Identification
(7) $M^{projected}_{occludee}=\text{ConcaveHull}(C).$
@@
(9) $M_{occlusion}=M_{union}\cap M_{occluder}.\vskip-3.00003pt$
With this approach, we precisely identify the occluded region of the target object without relying on any predefined templates, significantly enhancing the generalizability and robustness.
@@ 3.6. Application in 3D Reconstruction
(12) $\mathcal{L}_{obj}=\sum_{t}\mathcal{L}_{photo}\left(R(\mathcal{G}_{obj},H_{t}\circ X_{t}),\;\tilde{I}_{out}^{t}\right),\vskip-5.0pt$
@@ 4.3. Temporally Consistent Amodal Completion
We quantitatively evaluate our approach against several state-of-the-art baselines, including Pix2gestalt (Ozguroglu et al., 2024), Stable Diffusion Inpainting (Rombach et al., 2022), LaMa (Suvorov et al., 2022), and VDT (Lu et al., 2023), initialized with random noise. The results are summarized in Table 1, covering both human and object categories, as the human is treated as one of 19 object classes in our experiments, and only frames with an object occlusion ratio between 15% and 70% are considered. Our method consistently achieves the best performance across amodal completion metrics, IoU and CLIP, on both BEHAVE and InterCap. This indicates a higher fidelity in reconstructing occluded regions, validating the effectiveness of our template-free regioning strategy in complex HOI cases. To assess ideal condition performance, we report results using ground-truth masks in the last row.
@@
Figure 5. Qualitative results of 3D Reconstruction: Our method (left) produces clearer geometry and fewer artifacts than without our method (middle) or no completion (right).
@@
Table 2. Quantitative results on 3D reconstruction from the BEHAVE dataset (Bhatnagar et al., 2022): We report the metrics, computed within tight bounding boxes around humans or objects.
@@ C.1. Baselines
Table 5. 3D bounding box size and volume for each object in BEHAVE [4], sorted by descending size. boxtiny and toolbox have the smallest spatial extent and are excluded from 3D reconstruction experiments.
@@ D. Evaluation
To evaluate the quality of amodal completion, we use two complementary metrics: CLIP Score (Radford et al., 2021) and Intersection over Union (IoU). The CLIP score quantifies the semantic alignment between each inpainted image and a corresponding category-level textual prompt using CLIP’s vision–language embedding space. This metric has been widely adopted in various tasks (Saharia et al., 2022; Poole et al., 2022; Lee et al., 2025) without ground truth data to evaluate. This allows us to assess whether the completed region semantically resembles the intended object class. To reduce the impact of unrelated background pixels, we compute CLIP scores within tight bounding boxes centered on the reconstructed objects.
@@ E. Additional Experiments
Effect of Temporal Window Size.
F. Potential Applications
Our current work demonstrates novel pose synthesis by borrowing human and object poses from different videos. However, this approach can be significantly extended by integrating with text-to-motion and text-to-HOI models such as (Chi et al., 2024; Wu et al., 2024; Shi et al., 2025). By adapting human and object poses from the outputs of text prompts, our model can achieve greater flexibility in HOI animation.